OpenStudy (loser66):

how to write xy''-(cosx)y'+(sinx)y=0, x>0 under the form of [P(x)y']'+[f(x)y]'=0 Please, help

OpenStudy (tkhunny):

You could start from the back and give it a good look Using various rules (product, chain, etc) [P(x)y']' = P(x)y" + P'(x)y' [f(x)y]' = f(x)y' + f'(x)y'

OpenStudy (loser66):

[xy']' +[(-1-cos x)y]'=0 [xy']'= [(1+cos x)y]' integral both sides xy' =(1+cos x)y \[y' -\frac{1+cos x}{x}y=0\] \[\huge y = Ce^{-\int {-\frac{1+cos x}{x}}}\]

OpenStudy (loser66):

not that, ha, I got what I was wrong.

zepdrix (zepdrix):

oh u figured it out? :D

OpenStudy (loser66):

yes, thanks for respond. I got it.

OpenStudy (loser66):

\[y = e^{\mu t}\int Ce^\mu (t)dt +D\]

OpenStudy (loser66):

@ Chillout I am sorry friend. I figured out my mistake and carelessly wrote it down. it was not that. Actually it is \[y = \frac {1}{e^{\mu x dx}}\int C e^{\mu x dx} + D e^{-\ mu x dx}\] hopefully no more mistypo.

OpenStudy (anonymous):

Variation of parameters... what a fun mess! are you trying to solve the equation or just rewrite it bud?

OpenStudy (loser66):

@druminjosh I rewrote it first and then solved it. I knew how to solve it. At that moment, I didn't know how to link the forms. Now, I know it. thanks for respond.